# Maximal regularity for evolution equations with critical singular perturbations

### v1

@@ 1.2. Setting beyond maximal $L^{p}$-regularity
Th. 3.12 $X_{1}\to X_{\frac{1}{q},\infty}$ $\frac{1}{q}=\frac{1}{r}-\frac{1}{p}$ $L^{p}(I;X_{0})+L^{r}(I;X_{\frac{1}{q},\infty})$ $L^{p}(I;X_{1})\cap C(I;X_{\frac{1}{p^{\prime}},p})$
@@ 2.1. Maximal $L^{p}$-regularity
    $\|u\|_{\mathrm{MR}^{p}(0,b,w_{\kappa})}\leq\big(M(1\vee C_{A})+1\big)\|u_{0}\|_{X_{1-\frac{1+\kappa}{p},p}}+M\|f\|_{L^{p}(0,b,w_{\kappa};X_{0})}.$
@@
    $\|u\|_{\mathrm{MR}^{p}(a,b,w_{\kappa})}\leq T^{\kappa/p}\big(M(1\vee C_{A})+1\big)\|u_{a}\|_{X_{1-\frac{1}{p},p}}+M\|f\|_{L^{p}(a,b,w_{\kappa};X_{0})}.$
@@
applying (2.5) in the last line. Taking the infimum over all $z\in\mathrm{MR}^{p}(0,\infty,w_{\kappa})$, the result follows from the trace method for real interpolation, see the references above.
@@
Taking the infimum over all $z\in\mathrm{MR}^{p}(a,\infty)$, the result follows as before and by a simple shift. ∎
@@ 3.2. Embeddings of the mixed-scale maximal regularity space
The operator norm of the embedding depends on $T$.
@@
Next, we will prove trace embeddings. The embedding into $C_{\frac{\kappa}{p}}([0,T];X_{1-\frac{1}{p},p})$ plays a crucial role in the proof of our global well-posedness result, Theorem 3.12.
@@
$\hookrightarrow C([0,T];X_{1-\frac{1+\kappa}{p},r})\cap C_{\frac{\kappa}{p}}([0,T];X_{1-\frac{1}{p},r})$
$\hookrightarrow C([0,T];X_{1-\frac{1+\kappa}{p},p})\cap C_{\frac{\kappa}{p}}([0,T];X_{1-\frac{1}{p},p}).$
@@
$Z_{r}\hookrightarrow C_{\frac{\nu}{r}}([0,T];(X_{\gamma,\infty},X_{1+\gamma,\infty})_{1-\frac{1}{r},r})=C_{\frac{\nu}{r}}([0,T];X_{1+\gamma-\frac{1}{r},r}).$
@@
$\mathrm{MR}_{+}(0,T)$ $\hookrightarrow L^{p}(0,T,w_{\kappa};X_{1})\cap C([0,T];X_{1-\frac{1+\kappa}{p},p})\cap C_{\frac{\kappa}{p}}([0,T];X_{1-\frac{1}{p},p}).$
Note that the operator norm of the above embeddings depends on $T$ and blows up if $T\downarrow 0$ unless we assume that the functions vanish at $t=0$.
@@ 3.3. Main result
    For every $(p,\kappa)$-admissible triple $(r,\nu,\gamma)$ (see Definition 3.1), $A|_{X_{\gamma+1}}\colon[0,T]\to\mathcal{L}(X_{\gamma+1},X_{\gamma})$ is well-defined, strongly measurable in the strong operator topology, and $A|_{X_{\gamma+1}}$ has maximal $L^{r}(0,T,w_{\nu};X_{\gamma})$-regularity.
@@
In Assumption 3.7, $A|_{X_{\gamma+1}}(\cdot)x$ is defined for $x\in X_{1+\gamma}$, since $X_{\gamma+1}\hookrightarrow X_{\gamma+1,\infty}\hookrightarrow X_{1,1}\hookrightarrow X_{1}$. Moreover, Assumption 3.7(1) is satisfied if $X_{\gamma}$ is chosen as a complex interpolation space $X_{0},X_{1+\gamma_{*}}]_{\frac{\gamma}{1+\gamma_{*}}}$ or a real interpolation space $(X_{0},X_{1+\gamma_{*}})_{\frac{\gamma}{1+\gamma_{*}},q}$ with $q\in[1,\infty]$.
@@
In line with Definition 2.1, if $A$ satisfies Assumption 3.7, then we will write its maximal regularity constants as $M_{p,\kappa,A}(0,T)$ and $M_{r_{i},\nu_{i},A|_{X_{\gamma_{i}+1}}}(0,T)$, and we let
(3.16) $M_{T}\coloneqq M_{p,\kappa,A}(0,T)\vee\max_{1\leq i\leq k}M_{r_{i},\nu_{i},A|_{X_{\gamma_{i}+1}}}(0,T).$
@@
$p\in(1,\infty)$, $\kappa\geq 0$, $k\in\mathbb{N}$ and $(r_{i},\nu_{i},\gamma_{i})$ is $(p,\kappa)$-admissible for $i=1,\ldots,k$. We assume $B\colon[0,T]\to\mathcal{L}(X_{1},X_{0})$ decomposes as $B=\sum_{i=1}^{k}B_{i}$, where each $B_{i}\colon[0,T]\to\mathcal{L}(X_{1},X_{\gamma_{i}})$ is strongly measurable in the strong operator topology and
@@
In Assumption 3.9, $B_{i}=0$ whenever $p=r_{i}$. However, admissible triples $(r_{i},\nu_{i},\gamma_{i})$ with $p=r_{i}$ are retained to accommodate additional inhomogeneities $f_{i}$ later. Also, note that if $p>r_{i}$, then $(p,\kappa)$-admissibility implies that $\frac{\nu_{i}p-\kappa r_{i}}{p-r_{i}}\geq 0$ for the weight of $b_{i}$. Note that in applications to large deviation (see Remark 3.3), the weight in Assumption 3.9 becomes one. In this case $b_{i}\in L^{q_{i}}(0,T)$ with $\frac{1}{q_{i}}=\frac{1}{r_{i}}-\frac{1}{p}$.
@@
Moreover, since Assumption 3.9 fixes admissible triples, $\mathrm{MR}_{+}(0,T)$ is defined, and we will require the following mapping property of $A$, which is an analog of (2.5) in the current setting:
(3.18) $\begin{split}&\|Au\|_{L^{p}(0,T,w_{\kappa};X_{0})}\leq C_{A}\big(\|u\|_{L^{p}(0,T,w_{\kappa};X_{1})}+\|u\|_{C([0,T];X_{1-\frac{1+\kappa}{p},p})}+\|u\|_{C_{\nicefrac{{\kappa}}{{p}}}((0,T];X_{1-\frac{1}{p},p})}\big)\\ &\quad\text{ for all }u\in L^{p}(0,T,w_{\kappa};X_{1})\cap C([0,T];X_{1-\frac{1+\kappa}{p},p})\cap C_{\nicefrac{{\kappa}}{{p}}}((0,T];X_{1-\frac{1}{p},p}),\end{split}$
for a constant $C_{A}$.
Remark 3.11.
In (3.18), the function space for $u$ admits some flexibility: it suffices to estimate by $C_{A}\|u\|_{E(0,T)}$ for any normed space $E(0,T)$ such that $\mathrm{MR}_{+}(0,T)\hookrightarrow E(0,T)\hookrightarrow L^{p}(0,T,w_{\kappa};X_{1})$, as seen from the proof of Proposition 3.17. However, choosing the smallest space $E(0,T)=\mathrm{MR}_{+}(0,T)$ can lead to a constant $C_{A}$ that diverges as $T\downarrow 0$, which may be undesirable. Our current choice prevents this behavior and covers virtually all cases of interest.
@@
Let Assumptions 3.7 and 3.9 hold, let ${L_{+}}(0,T)$ and $\mathrm{MR}_{+}(0,T)$ be defined by (3.5) and (3.6), and let AA satisfy (3.18).
Let $u_{0}\in(X_{0},X_{1})_{1-\frac{1+\kappa}{p},p}$ and $f\in{L_{+}}(0,T)$. Then (3.3) has a unique strong solution $u\in\mathrm{MR}_{+}(0,T)$ and
(3.19) $\|u\|_{\mathrm{MR}_{+}(0,T)}+\|u\|_{L^{p}(0,T,w_{\kappa};X_{1})}$ $\leq C\Big(\|u_{0}\|_{X_{1-\frac{1+\kappa}{p},p}}+\|f\|_{{L_{+}}(0,T)}\Big),$
(3.20) $\|u\|_{C([0,T];X_{1-\frac{1+\kappa}{p},p})}+\|u\|_{C_{\frac{\kappa}{p}}([0,T];X_{1-\frac{1}{p},p})}$ $\leq C^{\prime}\Big(\|u_{0}\|_{X_{1-\frac{1+\kappa}{p},p}}+\|f\|_{{L_{+}}(0,T)}\Big),$
for constants $C$ and $C^{\prime}$ independent of $u_{0}$ and $f$.
The constant $C$ depends only on $b_{\scriptscriptstyle[0,T]}$, $M_{T}$ defined by (3.16), $C_{A}$ from (3.18), $p$, $r_{1},\ldots,r_{k}$, and $C^{\prime}=CK_{T}$, where $K_{T}$ is the norm of the embedding of Corollary 3.6.
@@
It readily follows from Theorem 3.12 that the claims therein remain true on subintervals $0,\tau]$ with $\tau\in(0,T]$, with the same constants $C$ and $C^{\prime}$ (i.e. dependent of $T$, not of $\tau$), even though the embedding constant of Corollary 3.6 blows up as $\tau\downarrow 0$.
Existence and uniqueness of solutions on $0,\tau]$ hold since Assumptions 3.7 and 3.9 are also satisfied on $0,\tau]$, due to Lemma 2.2, and (3.18) holds on $0,\tau]$ by applying a bounded extension operator to $u$. Concerning the constants $C$ and $C^{\prime}$, we can apply Theorem 3.12 to $f{{\bf 1}}_{[0,\tau]}\in{L_{+}}(0,T)$, restrict the corresponding solution (on $[0,T]$) to $0,\tau]$, and use uniqueness on $0,\tau]$.
When $A$ is a linear (time-independent) operator, the main theorem simplifies significantly, rendering several assumptions redundant.
@@
Put $\gamma_{*}\coloneqq 1$, $X_{2}\coloneqq D(A^{2})$ and write $X_{\gamma,q}\coloneqq(X_{0},D(A^{2}))_{\gamma/2,q}$ for $\gamma\in(0,2)$ and $q\in[1,\infty]$. Note that (3.18) holds since $X_{1}=D(A)$, so $\|Au\|_{L^{p}(0,T,w_{\kappa};X_{0})}\lesssim_{A,X_{1},X_{0}}\|u\|_{L^{p}(0,T,w_{\kappa};X_{1})}$ for any $u\in\mathrm{MR}_{+}(0,T)$. Hence, to prove the corollary, it remains to prove that Assumption 3.7 is satisfied for all three choices of $X_{\gamma}$. We can validate Assumption 3.7(2) at once: maximal $L^{p}(0,T;X_{0})$-regularity implies maximal $L^{p}(0,T,w_{\kappa};X_{0})$-regularity by [18, Prop. 17.2.36] (note that $\kappa/p\in[0,1-1/p)$ by Assumption 3.9). It remains to verify Assumption 3.7(1)(3). We do this case by case below. One more general observation will be useful: $(X_{0},D(A^{2}))_{\frac{1}{2},1}\hookrightarrow D(A)\hookrightarrow(X_{0},D(A^{2}))_{\frac{1}{2},\infty}$ by [29, §1.15.2 Th.(d)]. Therefore, using reiteration, we obtain
@@
Next, we return to the main theorem in which $A$ is time-dependent. Before we turn to the proof, we need several preparations. The next lemma summarizes the estimates that Assumption 3.7 gives for
@@
We use the notation of the maximal regularity constants from Definition 2.1.
@@
Let Assumption 3.7 hold and let $(r,\nu,\gamma)$ be $(p,\kappa)$-admissible. For all $\tau\in(0,T]$:
@@
  • For every $f\in L^{r}(0,\tau,w_{\nu};X_{\gamma})$, (3.27) has a unique strong solution $u\in L^{r}(0,\tau,w_{\nu};X_{1+\gamma})\cap W^{1,r}(0,\tau,w_{\nu};X_{\gamma})$, and
    $\|u\|_{L^{r}(0,\tau,w_{\nu};X_{1+\gamma})}+\|u\|_{W^{1,r}(0,\tau,w_{\nu};X_{\gamma})}\leq M_{r,\nu,A|_{X_{\gamma+1}}}(0,T)\|f\|_{L^{r}(0,\tau,w_{\nu};X_{\gamma})}.$

### v2

@@ 1.2. Setting beyond maximal $L^{p}$-regularity
Th. 3.11 $X_{1}\to X_{\frac{1}{q},\infty}$ $\frac{1}{q}=\frac{1}{r}-\frac{1}{p}$ $L^{p}(I;X_{0})+L^{r}(I;X_{\frac{1}{q},\infty})$ $L^{p}(I;X_{1})\cap C(I;X_{\frac{1}{p^{\prime}},p})$
@@ 2.1. Maximal $L^{p}$-regularity
    $\|u\|_{\mathrm{MR}^{p}(0,b,w_{\kappa})}\leq\tfrac{12p}{1+\kappa}\big(M(1\vee C_{A})+1\big)\|u_{0}\|_{X_{1-\frac{1+\kappa}{p},p}}+M\|f\|_{L^{p}(0,b,w_{\kappa};X_{0})}.$
@@
    $\|u\|_{\mathrm{MR}^{p}(a,b,w_{\kappa})}\leq 12pT^{\kappa/p}\big(M(1\vee C_{A})+1\big)\|u_{a}\|_{X_{1-\frac{1}{p},p}}+M\|f\|_{L^{p}(a,b,w_{\kappa};X_{0})}.$
@@
applying (2.5) in the last line. Taking the infimum over all $z\in\mathrm{MR}^{p}(0,\infty,w_{\kappa})$, the stated estimate follows from the trace method for real interpolation, see [18, Th. L.2.3] for the constant.
@@
Taking the infimum over all $z\in\mathrm{MR}^{p}(a,\infty)$, the claim follows as before and by a simple shift. ∎
@@ 3.2. Embeddings of the mixed-scale maximal regularity space
Let $p\in(1,\infty)$ and let $\kappa\in[0,p-1)$. Let $(r,\nu,\gamma)$ be $(p,\kappa)$-admissible. Then,
@@
The operator norm of the embedding depends on $p,\kappa,r,\nu,\gamma$, and $T$, and blows up if $T\downarrow 0$, unless we assume that the functions vanish at $t=0$.
@@
Next, we will prove trace embeddings. The embedding into $C_{\nicefrac{{\kappa}}{{p}}}((0,T];X_{1-\frac{1}{p},p})$ plays a crucial role in the proof of our global well-posedness result, Theorem 3.11.
@@
$\hookrightarrow C([0,T];X_{1-\frac{1+\kappa}{p},r})\cap C_{\nicefrac{{\kappa}}{{p}}}((0,T];X_{1-\frac{1}{p},r})$
$\hookrightarrow C([0,T];X_{1-\frac{1+\kappa}{p},p})\cap C_{\nicefrac{{\kappa}}{{p}}}((0,T];X_{1-\frac{1}{p},p}).$
@@
$Z_{r}\hookrightarrow C_{\nicefrac{{\nu}}{{r}}}((0,T];(X_{\gamma,\infty},X_{1+\gamma,\infty})_{1-\frac{1}{r},r})=C_{\nicefrac{{\nu}}{{r}}}((0,T];X_{1+\gamma-\frac{1}{r},r}).$
@@
$\mathrm{MR}_{+}(0,T)$ $\hookrightarrow L^{p}(0,T,w_{\kappa};X_{1})\cap C([0,T];X_{1-\frac{1+\kappa}{p},p})\cap C_{\nicefrac{{\kappa}}{{p}}}((0,T];X_{1-\frac{1}{p},p}).$
The operator norm of the embedding depends on $p,\kappa,r_{i},\nu_{i},\gamma_{i}$ for $1\leq i\leq k$, and $T$, and blows up if $T\downarrow 0$, unless we assume that the functions vanish at $t=0$.
@@ 3.3. Main result
    For every $(p,\kappa)$-admissible triple $(r,\nu,\gamma)$ (see Definition 3.1), there exists a map $A_{r,\nu,\gamma}\colon[0,T]\to\mathcal{L}(X_{\gamma+1},X_{\gamma})$ that is strongly measurable in the strong operator topology, and $A_{r,\nu,\gamma}$ has maximal $L^{r}(0,T,w_{\nu};X_{\gamma})$-regularity.
Moreover, the following mapping property holds for $A$ and $A_{r,\nu,\gamma}$, for every $(p,\kappa)$-admissible triple $(r,\nu,\gamma)$: there exist constants $C_{A}$ and $C_{A_{r,\nu,\gamma}}$ such that for all $u\in L^{p}(0,T,w_{\kappa};X_{1})\cap C([0,T];X_{1-\frac{1+\kappa}{p},p})\cap C_{\nicefrac{{\kappa}}{{p}}}((0,T];X_{1-\frac{1}{p},p})$, it holds that
(3.16) $\begin{cases}\|Au\|_{L^{p}(0,T,w_{\kappa};X_{0})}\leq C_{A}\Big(\|u\|_{L^{p}(0,T,w_{\kappa};X_{1})}+\|u\|_{C([0,T];X_{1-\frac{1+\kappa}{p},p})}+\|u\|_{C_{\nicefrac{{\kappa}}{{p}}}((0,T];X_{1-\frac{1}{p},p})}\Big)\\ \|A_{r,\nu,\gamma}u\|_{L^{p}(0,T,w_{\kappa};X_{0})}\leq C_{A_{r,\nu,\gamma}}\Big(\|u\|_{L^{p}(0,T,w_{\kappa};X_{1})}+\|u\|_{C([0,T];X_{1-\frac{1+\kappa}{p},p})}+\|u\|_{C_{\nicefrac{{\kappa}}{{p}}}((0,T];X_{1-\frac{1}{p},p})}\Big).\end{cases}$
We note that the mapping property (3.16) is similar to that in (2.5) used in Section 2. The version (3.16) will be sufficient to obtain uniqueness in the current setting. Other variants of this condition are also possible: see Remark 3.19.
@@
In Assumption 3.7(1), $X_{\gamma}$ can always be chosen as a complex interpolation space $X_{0},X_{1+\gamma_{*}}]_{\frac{\gamma}{1+\gamma_{*}}}$ or a real interpolation space $(X_{0},X_{1+\gamma_{*}})_{\frac{\gamma}{1+\gamma_{*}},q}$ with $q\in[1,\infty]$.
Moreover, the mappings $A_{r,\nu,\gamma}$ in Assumption 3.7(3) may be chosen freely, and need not be related to $A$. For instance $A_{r,\nu,\gamma}$ could be chosen time-independent. One could also choose $A_{r,\nu,\gamma}\coloneqq A|_{X_{\gamma+1}}$ (e.g. in the setting of Corollary 3.13). Note that $A|_{X_{\gamma+1}}(\cdot)x$ is defined for $x\in X_{1+\gamma}$, since $X_{\gamma+1}\hookrightarrow X_{\gamma+1,\infty}\hookrightarrow X_{1,1}\hookrightarrow X_{1}$.
@@
$p\in(1,\infty)$, $\kappa\geq 0$, $k\in\mathbb{N}$ and $(r_{i},\nu_{i},\gamma_{i})$ is $(p,\kappa)$-admissible for $i=1,\ldots,k$. Moreover, $B\colon[0,T]\to\mathcal{L}(X_{1},X_{0})$ decomposes as $B=\sum_{i=1}^{k}B_{i}$, where each $B_{i}\colon[0,T]\to\mathcal{L}(X_{1},X_{\gamma_{i}})$ is strongly measurable in the strong operator topology and
@@
In Assumption 3.9, $B_{i}=0$ whenever $p=r_{i}$. However, admissible triples $(r_{i},\nu_{i},\gamma_{i})$ with $p=r_{i}$ are retained to accommodate additional inhomogeneities $f_{i}$ later. Also, note that if $p>r_{i}$, then $(p,\kappa)$-admissibility implies that $\frac{\nu_{i}p-\kappa r_{i}}{p-r_{i}}\geq 0$ for the weight of $b_{i}$. Note that in applications to large deviations (see Remark 3.3), the weight in Assumption 3.9 becomes one. In this case, $b_{i}\in L^{q_{i}}(0,T)$ with $\frac{1}{q_{i}}=\frac{1}{r_{i}}-\frac{1}{p}$.
@@
Moreover, if $A$ and $A_{r,\nu,\gamma}$ are as in Assumption 3.7, then we will write their maximal regularity constants as $M_{p,\kappa,A}(0,T)$ and $M_{r,\nu,A_{r,\nu,\gamma}}(0,T)$, in line with Definition 2.1. Moreover, under Assumption 3.9 (so that admissible triples are fixed), we let
(3.18) $M_{T}\coloneqq M_{p,\kappa,A}(0,T)\vee\max_{1\leq i\leq k}M_{r_{i},\nu_{i},A_{r_{i},\nu_{i},\gamma_{i}}}(0,T).$
@@
Let Assumptions 3.7 and 3.9 hold. Let ${L_{+}}(0,T)$ and $\mathrm{MR}_{+}(0,T)$ be defined by (3.5) and (3.6). Let $u_{0}\in(X_{0},X_{1})_{1-\frac{1+\kappa}{p},p}$ and $f\in{L_{+}}(0,T)$.
Then, (3.3) has a unique strong solution $u\in\mathrm{MR}_{+}(0,T)$ and
$\begin{split}\|u\|_{\mathrm{MR}_{+}(0,T)}+&\|u\|_{L^{p}(0,T,w_{\kappa};X_{1})}+\|u\|_{C([0,T];X_{1-\frac{1+\kappa}{p},p})}+\|u\|_{C_{\nicefrac{{\kappa}}{{p}}}((0,T];X_{1-\frac{1}{p},p})}\\ &\leq C\big(\|u_{0}\|_{X_{1-\frac{1+\kappa}{p},p}}+\|f\|_{{L_{+}}(0,T)}\big),\end{split}$
for a constant $C$ independent of $u_{0}$ and $f$. The constant $C$ depends only on $p$, $\kappa$, and $r_{i},\nu_{i},\gamma_{i}$ for $1\leq i\leq k$, and nondecreasingly on $b_{\scriptscriptstyle[0,T]}$, $M_{T}$ defined by (3.18), and $C_{A}$ and $C_{A_{r_{i},\nu_{i},\gamma_{i}}}$ from (3.16).
@@
It readily follows from Theorem 3.11 that the claims therein remain true on subintervals $0,\tau]$ with $\tau\in(0,T]$, with the same constant $C$ (i.e. for the case of the full interval $[0,T]$).
Existence and uniqueness of solutions on $0,\tau]$ hold since Assumptions 3.7 and 3.9 are also satisfied on $0,\tau]$, due to Lemma 2.2 and since (3.16) holds on $0,\tau]$ (with new constants) by applying a bounded extension operator to $u$.
Concerning the final constant $C$, we can apply Theorem 3.11 to $f{{\bf 1}}_{[0,\tau]}\in{L_{+}}(0,T)$, restrict the corresponding solution (on $[0,T]$) to $0,\tau]$, and use uniqueness on $0,\tau]$.
When $A$ is a (time-independent) linear operator, the main theorem simplifies significantly, rendering several assumptions redundant.
@@
Put $\gamma_{*}\coloneqq 1$, $X_{2}\coloneqq D(A^{2})$ and write $X_{\gamma,q}\coloneqq(X_{0},D(A^{2}))_{\gamma/2,q}$ for $\gamma\in(0,2)$ and $q\in[1,\infty]$. Define $A_{r,\nu,\gamma}\coloneqq A|_{X_{\gamma+1}}$ whenever $(r,\nu,\gamma)$ is a $(p,\kappa)$-admissible triple.
To prove the corollary, it remains to prove that Assumption 3.7 is satisfied for all three choices of $X_{\gamma}$, and with $A_{r,\nu,\gamma}$ defined above. Note that (3.16) holds since $X_{1}=D(A)$, so $\|Au\|_{L^{p}(0,T,w_{\kappa};X_{0})}\lesssim_{A,X_{1},X_{0}}\|u\|_{L^{p}(0,T,w_{\kappa};X_{1})}$ for any $u\in L^{p}(0,T,w_{\kappa};X_{1})$. Moreover, we can validate Assumption 3.7(2) at once: maximal $L^{p}(0,T;X_{0})$-regularity implies maximal $L^{p}(0,T,w_{\kappa};X_{0})$-regularity by [18, Prop. 17.2.36] (note that $\kappa/p\in[0,1-1/p)$ by Assumption 3.9). It remains to verify Assumption 3.7(1)(3). We do this case by case below. One more general observation will be useful: $(X_{0},D(A^{2}))_{\frac{1}{2},1}\hookrightarrow D(A)\hookrightarrow(X_{0},D(A^{2}))_{\frac{1}{2},\infty}$ by [29, §1.15.2 Th.(d)]. Therefore, using reiteration, we obtain
@@
We return to the main theorem, Theorem 3.11, in which $A$ is time-dependent. Before proceeding to the proof, we need several preparations. The next lemma establishes estimates that Assumption 3.7 gives for
@@
We use the notation of the maximal regularity constants from Definition 2.1. Note that uniqueness for part (ii) of the lemma below will be proved more generally in Proposition 3.17.
@@
Let Assumption 3.7 hold and let $(r,\nu,\gamma)$ be $(p,\kappa)$-admissible. Then, for all $\tau\in(0,T]$, we have:
  • (i)
@@
  • (ii)
    For every $f\in L^{r}(0,\tau,w_{\nu};X_{\gamma})$, (3.25) has a strong solution $u\in F(0,\tau)$ such that
    $\|u\|_{F(0,\tau)}\leq M_{r,\nu,A_{r,\nu,\gamma}}(0,T)(1+M_{p,\kappa,A}(0,T)L(C_{A_{r,\nu,\gamma}}+C_{A}))\|f\|_{L^{r}(0,\tau,w_{\nu};X_{\gamma})},$
    where $L$ is a constant depending only on $p,\kappa,r,\nu,$ and $\gamma$, and $C_{A}$ and $C_{A_{r,\nu,\gamma}}$ are the constants from (3.16), and
    $F(0,\tau)\coloneqq\big(L^{p}(0,\tau,w_{\kappa};X_{1})\cap W^{1,p}(0,\tau,w_{\kappa};X_{0})\big)+\big(L^{r}(0,\tau,w_{\nu};X_{\gamma+1})\cap W^{1,r}(0,\tau,w_{\nu};X_{\gamma})\big).$